# A Bayesian Nonparametric Stochastic Block Model for Directed Acyclic Graphs

### v2

@@ 2.1 Likelihood
With all required quantities defined, we will derive the likelihood with observed $\mathbf{Y}$ given $\mathbf{Z}_{n}$ and $\boldsymbol{\sigma}$. We first check that $\boldsymbol{\sigma}$ is topological, or equivalently $\mathbf{Y}^{\boldsymbol{\sigma}}$ is upper triangular i.e. $Y^{\boldsymbol{\sigma}}_{qp}=0$ for all dyads $(p,q)$ where $1\leq p<q\leq n$, otherwise the likelihood is 0. Once $\boldsymbol{\sigma}$ is checked to be topological, using Equation 2, the observed data likelihood is
@@
where $\bar{Y}=\prod_{p=1}^{n-1}\prod_{q=p+1}^{n}\left(Y^{\boldsymbol{\sigma}}_{pq}!\right)^{-1}$, ${}\displaystyle M_{ij}=\sum_{p=1}^{n-1}\sum_{q=p+1}^{n}\xi^{\boldsymbol{\sigma}}_{p}\xi^{\boldsymbol{\sigma}}_{q}\mathbb{I}(Z^{\boldsymbol{\sigma}}_{p}=i,Z^{\boldsymbol{\sigma}}_{q}=j)$, and ${}\displaystyle E_{ij}=\sum_{p=1}^{n-1}\sum_{q=p+1}^{n}Y^{\boldsymbol{\sigma}}_{pq}\mathbb{I}(Z^{\boldsymbol{\sigma}}_{p}=i,Z^{\boldsymbol{\sigma}}_{q}=j)$. Detailed derivations are in the online Appendix B. Likelihood 3 is influenced by $\mathbf{Z}_{n}$ and $\boldsymbol{\sigma}$ through the two matrices $\mathbf{E}$ and $\mathbf{M}$.
@@ 2.2 Priors and posterior density
For $\mathbf{C}$, we assume each $C_{ij}$ is a priori independent and identically distributed according to the Gamma$(a,b)$ distribution, where $a$ and $b$ are the positive shape and rate parameters, respectively. This enables $\mathbf{C}$ to be integrated out, to obtain
[…]
$\quad=\mathbb{I}(\mathbf{Y}^{\boldsymbol{\sigma}}\text{ upper tri.})\times\bar{Y}\times\prod_{p=1}^{n}\prod_{q=1}^{n}\left(\xi_{p}\xi_{q}\right)^{Y_{pq}}\times\left(\frac{b^{a}}{\Gamma(a)}\right)^{K_{n}^{2}}\prod_{i=1}^{K_{n}}\prod_{j=1}^{K_{n}}\frac{\Gamma(E_{ij}+a)}{(M_{ij}+b)^{E_{ij}+a}}$ (4)
@@
For $\boldsymbol{\sigma}$, we assign a uniform prior to all permutations of $\{1,2,\ldots,n\}$ i.e. $\pi(\boldsymbol{\sigma})=(n!)^{-1}$. There is no issue with an ordering that is not topological having a positive prior probability, as such an ordering will result in $\mathbb{I}(\mathbf{Y}^{\boldsymbol{\sigma}}\text{ upper tri.})$. ) and the likelihood (4) being equal to 0.
@@
For $\mathbf{Z}_{n}$, we assume that it comes from an exchangeable sequence $(Z_{p})_{p\in\mathbb{N}}$ driven by a Pitman-Yor process, Pitman and Yor (1997), i.e. let $Z_{p}|P\stackrel{{\scriptstyle iid}}{{\sim}}P$ for all $p\in\mathbb{N}$ and $P\sim\text{PY}(\alpha,\theta,P_{0})$. The Pitman-Yor process, reviewed in the online Appendix A.3, is a distribution for an unknown probability distribution $P$ and is parametrized by three hyperparameters $(\alpha,\theta,P_{0})$, where $P_{0}$, called base distribution, is a distribution on the sample space, and $\alpha$ and $\theta$ are two scalars satisfying either: 1) $10\leq\alpha<$ 1 and $\theta\geq-\alpha$; 2) $\alpha<0$ and $\theta=k|\alpha|$ for $k\in\mathbb{N}$. Case 1) corresponds to the infinite regime, in which $P$ has an infinite number of support points, and the sequence $(Z_{p})_{p\in\mathbb{N}}$ will display infinite number of distinct values. Case 2) is the finite regime and $P$ has $k$ support points, which is also the total number of distinct values in $(Z_{p})_{p\in\mathbb{N}}$. The Dirichlet process corresponds to the special case $\alpha=0$. When considering the finite regime, i.e. when $\alpha<0$, we apply the re-parametrization $(\alpha,\theta)\to(\gamma,k)$, with $\gamma:=|\alpha|>0$ and $k:=\theta/|\alpha|\in\mathbb{N}$, and assign a prior to $(\gamma,k)\in\mathbb{R}_{+}\times\mathbb{N}$.
@@
As detailed in the online Appendix A.3, it is possible to compute the marginal likelihood (when $P$ is marginalized out) of a sample $\mathbf{Z}_{n}$ driven by the PY process, which we denoted by $\mathbb{P}(\mathbf{Z}_{n}|\boldsymbol{\eta}_{r})$, where $\boldsymbol{\eta}_{r}$ is a parameter vector of length 2, dependent on the choice of regime $r\in\{0,1\}$, either infinite $\boldsymbol{\eta}_{0}=(\alpha,\theta)$ or finite $\boldsymbol{\eta}_{1}=(\gamma,k)$. This in turn requires the specification of the prior of $\boldsymbol{\eta}_{r}$ under both regimes. Under the infinite regime, $r=0$ and $\boldsymbol{\eta}_{r}=\boldsymbol{\eta}_{0}=(\alpha,\theta)$, and we assume that $\alpha\sim\text{Uniform}[0,1]$ and $\theta+\alpha$ follows a Gamma distribution. This choice of dependent priors for $\alpha$ and $\theta$ is done in order to include all possible values in parameter space. Indeed, when $10<\alpha<$ 1, the range of possible values of $\theta$ is $\theta>-\alpha$. Under the finite regime, $r=1$ and $\boldsymbol{\eta}_{r}=\boldsymbol{\eta}_{1}=(\gamma,k)$, and we assume that $\gamma$ and $k$ are independent a priori, $\gamma$ follows a Gamma distribution, and $k$ follows a truncated negative binomial distribution, with parameters $a_{k}$ and $b_{k}$, and density
@@
where $\Gamma(\cdot)$ is the gamma function. The factor $(1-b_{k}^{a_{k}})^{-1}$ is due to the truncation of 00 from the original support of the negative binomial distribution. The negative binomial is over-dispersed, allowing possibly high variance, hence high prior uncertainty of the value of $k$.
@@
The regime-dependent parameters and their priors are introduced this way because, ultimately, we want to enable model selection of the regime, which in turn requires the prior of $r$, denoted by $\mathbb{P}(r)$. The boundary cases $\mathbb{P}(r=0)=1$ and $\mathbb{P}(r=1)=1$ represent staying within the infinite and finite regimes, respectively, while the incorporation of the model selection in the MCMC sampler allows the data to select the more suitable among the two regimes. The model selection step, implemented within the MCMC sampler, follows the algorithm of Carlin and Chib (1995). Moreover, the prior choice for $r$ allows unbalanced weights between models. This might improve mixing of the model selection step, as shown in Friel and Pettitt (2008), in the context of reversible jump MCMC. At the end of Section 3.1, in Table 1, we show results from a simulated study in which the posterior sampler seems to be able to recover the correct regime between the finite and infinite ones. A detailed description of the model selection step can be found in the online Appendix C.2.
@@
As all priors required have been specified, the joint posterior of $\mathbf{Z}_{n}$, $\boldsymbol{\sigma}$, $\boldsymbol{\eta}_{r}$, $a$ and $b$ (and $r$), up to a proportionality constant, is readily available:
$\mathbb{P}(\mathbf{Z}_{n},\boldsymbol{\sigma},\boldsymbol{\xi},\boldsymbol{\eta}_{r},a,b,r|Y)\propto\mathbb{P}(\mathbf{Y}|\mathbf{Z}_{n},\boldsymbol{\sigma},\boldsymbol{\xi},a,b)\mathbb{P}(\mathbf{Z}_{n}|\boldsymbol{\eta}_{r})\mathbb{P}(\boldsymbol{\sigma})\mathbb{P}(\boldsymbol{\xi})\mathbb{P}(\boldsymbol{\eta}_{r}|r)\mathbb{P}(a)\mathbb{P}(b)\mathbb{P}(r).$ (5)
@@
We carry out Bayesian inference by sampling from this joint posterior using MCMC. The full detail description of the MCMC sampler is available in the online Appendix C.1.
@@ 3.1 Simulated data
This section presents an illustrative simulation study in which the proposed DAG-SBM (Equation 2), is compared with the SBM for a directed graph (Equation 1, referred to as the directed SBM). The purpose of this study is to show how fitting the directed SBM without considering the inherent hierarchical ordering among vertices can result in assigning positive posterior probability to many “impossible” edges, and substantial bias in the estimate of the overall number of blocks in the model.
@@
Upon sampling the parameters from their priors specified in Section 2.2, adjacency matrices are sampled from the DAG-SBM for different sample sizes $n=\{250$,$500$,1000}1000\}1000 } and combinations of the Pitman-Yor parameters, $(\alpha,\theta)=\{(-1,10)$, $(0,1)$, $(0.1,1)$, (0.6,1)}(0.6,1)\}( 0.6 , 1 ) }, thus including both finite and infinite regimes. Both the DAG-SBM and directed SBM are fitted to the resulting adjacency matrices, also under both the finite and infinite regimes.
@@
Figure 1: Actual adjacency matrix (left) for data simulated from the DAG-SBM (Equation 2) with $n=250$ and different values of $(\alpha,\theta)$, and estimated probabilities by the directed SBM (Equation 1) under the finite (middle column) and infinite (right) regime, ordered by true $\boldsymbol{\sigma}$. The middle and right plots correspond to the results in finite and infinite regimes in the corresponding bottom left panels of Figure 2. The dashed line is the major diagonal.
@@
Figure 1 plots the heat maps of the actual adjacency matrix $\mathbf{Y}^{\boldsymbol{\sigma}}$ with $n=250$ (not shown for $n=500$ and $n=1000$), ordered according to the true topological ordering, alongside with the posterior point estimates of $\mathbf{Y}^{\boldsymbol{\sigma}}$, for each simulated data set fit by the directed SBM. The true adjacency matrices in the first column are necessarily upper triangular, given that they have been reordered according to the true $\boldsymbol{\sigma}$. However, from the second and third columns, it is evident that many edges in the lower triangular sections have positive posterior probability, while it should be equal to zero. This is especially true for vertices having either high in-degree or out-degree counts. This is due to the fact that the directed SBM is row and column exchangeable and vertices with high in-degree tends to have also high out-degree counts. However, when a hierarchical ordering is present, the model is not exchangeable anymore and the out- and in-degrees of each vertex should be negatively correlated random variables, with vertices at the beginning the topological order having higher out-degree than in-degree counts, while the opposite for vertices at the end of the topological order.
Figures 2 displays the posterior histograms of the number of blocks $K_{n}$ estimated using the DAG-SBM and the directed SBM, for the same sets of simulated data. On the one hand, the directed SBM, which does not include the topological ordering in the likelihood, has the tendency of heavily underestimating or overestimating $K_{n}$. The increase in the number of blocks also negatively affects the computational costs of the algorithm, compared to the DAG version. A comparison of the computational times between the two models is provided in the online Appendix D.1. On the other hand, the DAG-SBM, which accounts for the topological ordering, provides reasonable estimates of $K_{n}$ across different sample sizes and for both regimes of the Pitman-Yor parameters. Furthermore, for the DAG-SBM, the model selection step in the MCMC sampler successfully recovers the true regime of the Pitman-Yor parameters. Indeed, Table 1 displays the estimated Bayes factor for the model selection step between the finite and infinite regime, for a range of different values of the hyperparameters of the Pitman-Yor process. The estimated Bayes factor suggests that the posterior sampler is capable at recovering the correct regime from the data, particularly when $\alpha$ is far from 00, and being generally indifferent between the regimes when $\alpha$ is close to 00.
@@
Figure 2: Posterior histogram of $K_{n}$ for data simulated from the DAG-SBM (Equation 2), with different values of $(\alpha,\theta,n)$, fitted by both the DAG-SBM (top row) and directed SBM (Equation 1, bottom row), under the finite (solid) and infinite (stripe) regimes. The vertical dashed line is the true value of $K_{n}$.
$\alpha$ $\theta$ $n=250$ $n=500$ $n=1000$
$-1$ $10$ $8.30$ $612.58$ $2.72$
00 $1$ $2.56$ $0.85$ $19.12$
$0.1$ $1$ $1.11$ $0.40$ $0.89$
$0.6$ $1$ $7.93\times 10^{-4}$ $7.62\times 10^{-7}$ $1.92\times 10^{-11}$
@@
Table 1: Estimated Bayes factor for the finite regime to the infinite regime when applying the model selection in the MCMC sampler for the DAG-SBM to the simulated data, for different values of $(\alpha,\theta,n)$.
@@
Finally, we look at the posterior of $\boldsymbol{\sigma}$, or equivalently the positions of the vertices in the topological ordering, $\boldsymbol{\phi}$. The mixing of the MCMC is sufficiently good that the trace plots are not shown. The posterior density of each component of $\boldsymbol{\phi}$ is plotted as a row in Figure 3, with the rows in the true topological ordering. The alignment suggests that this latent ordering can be well recovered without identifiability issues.
@@
Figure 3: The posterior density of the positions of the vertices in $\boldsymbol{\sigma}$ for $n=250$ and different values of $(\alpha,\theta)$ for data simulated from the DAG-SBM (Equation 2), fitted by the DAG-SBM under the true regime according to $\alpha$. Each row of colours is the posterior density of a component of $\boldsymbol{\phi}$.

### v3

@@ 2.1 Likelihood
With the model established in Equation 2, we can proceed to compute the likelihood. As $\mathbf{Z}_{n}$ is unknown prior to fitting the model, it will be treated as a vector of latent variables and assigned a prior, of which the parametrisation will be detailed in Section 2.2. However, given this prior choice, several quantities can be defined to facilitate the likelihood derivations. First, there will be $K_{n}$ distinct values $(Z_{1}^{*},\ldots,Z_{K_{n}}^{*})$ appearing in $\mathbf{Z}_{n}$ with respective frequencies $\mathbf{N}:=(N_{1},\ldots,N_{K_{n}})$. The $i$-th element of $\mathbf{N}$, i.e. $N_{i}=\sum_{p=1}^{n}\mathbb{I}(Z_{p}=Z_{i}^{*})=\sum_{p=1}^{n}\mathbb{I}(Z_{p}=i)$ is then the number of vertices in group $i$. Second, we derive from $\mathbf{Z}_{n}$ and $\mathbf{Y}$ (or $\mathbf{Y}^{\boldsymbol{\sigma}}$) two $K_{n}\times K_{n}$ matrices $\mathbf{E}:=(E_{ij})_{1\leq{}i,j\leq{}K_{n}}$ and $\mathbf{M}:=(M_{ij})_{1\leq{}i,j\leq{}K_{n}}$. The matrix $\mathbf{E}$ is the edge matrix between the groups, where
$E_{ij}=\sum_{p=1}^{n}\sum_{q=1}^{n}Y_{pq}\mathbb{I}(Z_{p}=i,Z_{q}=j)=\sum_{p=1}^{n-1}\sum_{q=p+1}^{n}Y^{\boldsymbol{\sigma}}_{pq}\mathbb{I}(Z^{\boldsymbol{\sigma}}_{p}=i,Z^{\boldsymbol{\sigma}}_{q}=j),$ (3)
while in the matrix $\mathbf{M}$, $M_{ij}$ is the number of dyads $(p,q)$ between groups $i$ and $j$ such that vertex $p$ is topologically in front of vertex $q$. Mathematically,
$M_{ij}=\sum_{p=1}^{n-1}\sum_{q=p+1}^{n}\xi^{\boldsymbol{\sigma}}_{p}\xi^{\boldsymbol{\sigma}}_{q}\mathbb{I}\left(Z_{\sigma_{p}}=i,Z_{\sigma_{q}}=j\right)=\sum_{p=1}^{n-1}\sum_{q=p+1}^{n}\xi^{\boldsymbol{\sigma}}_{p}\xi^{\boldsymbol{\sigma}}_{q}\mathbb{I}\left(Z^{\boldsymbol{\sigma}}_{p}=i,Z^{\boldsymbol{\sigma}}_{q}=j\right).$ (4)
With all required quantities defined, we will derive the likelihood with observed $\mathbf{Y}$ given $\mathbf{Z}_{n}$ and $\boldsymbol{\sigma}$. We first check that $\boldsymbol{\sigma}$ is topological, or equivalently $\mathbf{Y}^{\boldsymbol{\sigma}}$ is upper triangular, i.e. $Y^{\boldsymbol{\sigma}}_{qp}=0$ for all dyads $(p,q)$ where $1\leq p<q\leq n$, otherwise the likelihood is 0. Once $\boldsymbol{\sigma}$ is checked to be topological, using Equation (2), the observed data likelihood is
@@
where $\bar{Y}=\prod_{p=1}^{n-1}\prod_{q=p+1}^{n}\left(Y^{\boldsymbol{\sigma}}_{pq}!\right)^{-1}$, and $E_{ij}$ and $M_{ij}$ are given by Equations 3 and 4, respectively. Detailed derivations are in the online Appendix B. Likelihood (5) is influenced by $\mathbf{Z}_{n}$ and $\boldsymbol{\sigma}$ through the two matrices $\mathbf{E}$ and $\mathbf{M}$.
@@ 2.2 Priors and posterior density
For $\mathbf{C}$, we assume each $C_{ij}$ is a priori independent and identically distributed according to the Gamma(a,b)(a,b) distribution, where $a$ and $b$ are the positive shape and rate parameters, respectively. This enables $\mathbf{C}$ to be integrated out, to obtain
$\mathbb{P}(\mathbf{Y}|\mathbf{Z}_{n},\boldsymbol{\sigma},\boldsymbol{\xi},a,b)=\int\mathbb{P}(\mathbf{Y}|\mathbf{C},\mathbf{Z}_{n},\boldsymbol{\sigma},\boldsymbol{\xi})\mathbb{P}(\mathbf{C}|a,b)d\mathbf{C}$
$\quad=\mathbb{I}(\mathbf{Y}^{\boldsymbol{\sigma}}\text{ upper tri.})\times\bar{Y}\times\prod_{p=1}^{n}\prod_{q=1}^{n}\left(\xi_{p}\xi_{q}\right)^{Y_{pq}}\times\left(\frac{b^{a}}{\Gamma(a)}\right)^{K_{n}^{2}}\prod_{i=1}^{K_{n}}\prod_{j=1}^{K_{n}}\frac{\Gamma(E_{ij}+a)}{(M_{ij}+b)^{E_{ij}+a}}.$ (6)
@@
For $\boldsymbol{\sigma}$, we assign a uniform prior to all permutations of $\{1,2,\ldots,n\}$, i.e. $\pi(\boldsymbol{\sigma})=(n!)^{-1}$. There is no issue with an ordering that is not topological having a positive prior probability, as such an ordering will result in $\mathbb{I}(\mathbf{Y}^{\boldsymbol{\sigma}}\text{ upper tri.})$ and the likelihood (6) being equal to 0.
@@
For $\mathbf{Z}_{n}$, we assume that its components are the first $n$ elements of an exchangeable sequence $(Z_{p})_{p\in\mathbb{N}}$ driven by a Pitman-Yor (PY) process (Pitman and Yor, 1997), i.e. $Z_{p}|P\stackrel{{\scriptstyle iid}}{{\sim}}P$ for all $p\in\mathbb{N}$ and $P\sim\text{PY}(\alpha,\theta,P_{0})$. The PY process, reviewed in the online Appendix A.3, is a distribution for an unknown probability distribution $P$ and is parametrised by three hyperparameters $(\alpha,\theta,P_{0})$, where $P_{0}$, called base distribution, is a distribution on the sample space, and $\alpha$ and $\theta$ are two scalars satisfying either: 1) $0\leq\alpha<1$ and $\theta\geq-\alpha$; 2) $\alpha<0$ and $\theta=k|\alpha|$ for $k\in\mathbb{N}$. The Dirichlet process corresponds to the special case $\alpha=0$.
2.3 Bayesian model selection
The PY process prior for $\mathbf{Z}_{n}$ helps select the number of blocks and distinguish between the infinite and finite regimes. Specifically, in the infinite regime, corresponding to $0\leq\alpha<1$ and $\theta\geq-\alpha$, $P$ has an infinite number of support points, and the sequence $(Z_{p})_{p\in\mathbb{N}}$ will display an infinite number of distinct values. In the finite regime, corresponding to $\alpha<0$ and $\theta=k|\alpha|$ for $k\in\mathbb{N}$, $P$ has $k$ support points, which is also the total number of distinct values in $(Z_{p})_{p\in\mathbb{N}}$. When considering the finite regime, i.e. when $\alpha<0$, we apply the re-parametrisation $(\alpha,\theta)\to(\gamma,k)$, with $\gamma:=|\alpha|>0$ and $k:=\theta/|\alpha|\in\mathbb{N}$, and assign a prior to $(\gamma,k)\in\mathbb{R}_{+}\times\mathbb{N}$.
@@ 2.3 Bayesian model selection
As detailed in the online Appendix A.3, it is possible to compute the marginal likelihood (when $P$ is marginalised out) of a sample $\mathbf{Z}_{n}$ driven by the PY process, which we denoted by $\mathbb{P}(\mathbf{Z}_{n}|\boldsymbol{\eta}_{r})$, where $\boldsymbol{\eta}_{r}$ is a parameter vector of length 2, dependent on the choice of regime $r\in\{0,1\}$, either infinite $\boldsymbol{\eta}_{0}=(\alpha,\theta)$ or finite $\boldsymbol{\eta}_{1}=(\gamma,k)$. This in turn requires the specification of the prior of $\boldsymbol{\eta}_{r}$ under both regimes. Under the infinite regime, $r=0$ and $\boldsymbol{\eta}_{r}=\boldsymbol{\eta}_{0}=(\alpha,\theta)$, and we assume that $\alpha\sim\text{Uniform}[0,1]$ and $\theta+\alpha$ follows a Gamma distribution. This choice of dependent priors for $\alpha$ and $\theta$ is done in order to include all possible values in parameter space. Under the finite regime, $r=1$ and $\boldsymbol{\eta}_{r}=\boldsymbol{\eta}_{1}=(\gamma,k)$, and we assume that $\gamma$ and $k$ are independent a priori, $\gamma$ follows a Gamma distribution, and $k$ follows a truncated negative binomial distribution, with parameters $a_{k}$ and $b_{k}$, and density
@@
where $\Gamma(\cdot)$ is the gamma function. The factor $(1-b_{k}^{a_{k}})^{-1}$ is due to the truncation of 0 from the original support of the negative binomial distribution. The negative binomial is over-dispersed, allowing possibly high variance, hence high prior uncertainty of the value of $k$.
@@
The regime-dependent parameters and their priors are introduced this way because, ultimately, we want to enable model selection of the regime, which in turn requires the prior of $r$, denoted by $\mathbb{P}(r)$. The boundary cases $\mathbb{P}(r=0)=1$ and $\mathbb{P}(r=1)=1$ represent staying within the infinite and finite regimes, respectively, while model selection takes place when $0<\mathbb{P}(r=0)<1$, allowing the data to select the more suitable among the two regimes. The model selection step, implemented within the MCMC sampler, follows the algorithm of Carlin and Chib (1995). Moreover, the prior choice for $r$ allows unbalanced weights between models, as this may improve mixing of the model selection step, as shown in Friel and Pettitt (2008), in the context of reversible jump MCMC. At the end of Section 3.1, in Table 1, we show results from a simulated study in which the posterior sampler in general recovers the correct regime between the finite and infinite ones. A detailed description of the model selection step can be found in the online Appendix C.2.
@@
As all priors required have been specified, the joint posterior of $\mathbf{Z}_{n}$, $\boldsymbol{\sigma}$, $\boldsymbol{\eta}_{r}$, $a$ and $b$ (and $r$), up to a proportionality constant, is
$\mathbb{P}(\mathbf{Z}_{n},\boldsymbol{\sigma},\boldsymbol{\xi},\boldsymbol{\eta}_{r},a,b,r|\mathbf{Y})\propto\mathbb{P}(\mathbf{Y}|\mathbf{Z}_{n},\boldsymbol{\sigma},\boldsymbol{\xi},a,b)\mathbb{P}(\mathbf{Z}_{n}|\boldsymbol{\eta}_{r})\mathbb{P}(\boldsymbol{\sigma})\mathbb{P}(\boldsymbol{\xi})\mathbb{P}(\boldsymbol{\eta}_{r}|r)\mathbb{P}(a)\mathbb{P}(b)\mathbb{P}(r).$ (7)
@@
We sample from this joint posterior using MCMC, of which the detailed description is available in the online Appendix C.1.
2.4 Posterior point estimate of $\mathbf{Z}_{n}$
To provide a posterior point estimate of $\mathbf{Z}_{n}$, we follow the clustering approach introduced in Meilă (2007) and further discussed in Wade and Ghahramani (2018). Specifically, the point estimate, denoted by $\hat{\mathbf{Z}}_{n}$, is obtained using a decision theoretic approach, by minimizing with respect to the posterior distribution a loss function on the space of allocation vectors,
For the loss function $L(\mathbf{Z}_{n},\tilde{\mathbf{Z}}_{n})$, Meilă (2007) chose the Variation of Information (VI), defined as
$\text{VI}(\mathbf{Z}_{n},\tilde{\mathbf{Z}}_{n})=\sum_{i=1}^{K_{n}}\frac{n_{i+}}{n}\log\left(\frac{n_{i+}}{n}\right)+\sum_{j=1}^{\tilde{K}_{n}}\frac{n_{+j}}{n}\log\left(\frac{n_{+j}}{n}\right)-2\sum_{i=1}^{K_{n}}\sum_{j=1}^{\tilde{K}_{n}}\frac{n_{ij}}{n}\log\left(\frac{n_{ij}}{n}\right),$
where $n_{ij}=\sum_{p=1}^{n}\mathbb{I}(Z_{p}=i,\tilde{Z}_{p}=j)$, $n_{i+}=\sum_{j=1}^{\tilde{K}_{n}}n_{ij}$, and $n_{+j}=\sum_{j=1}^{K_{n}}n_{ij}$. The loss function $L(\mathbf{Z}_{n},\tilde{\mathbf{Z}}_{n})$ can be seen as a distance between $\mathbf{Z}_{n}$ and $\tilde{\mathbf{Z}}_{n}$, which can be computed even if $K_{n}\neq\tilde{K}_{n}$, i.e. the numbers of groups implied by $\mathbf{Z}_{n}$ and $\tilde{\mathbf{Z}}_{n}$ are different.
2.5 Identifiability of $\mathbf{Z}_{n}$ and $\boldsymbol{\sigma}$
Our inference approach does not raise any concern of identifiability issues. For $\mathbf{Z}_{n}$, the problem of label switching in Bayesian MCMC estimation of finite mixture models and related models is indeed less common in nonparametric models, given that the prior of the block weights is not symmetric (Jasra et al., 2005, Papaspiliopoulos and Roberts, 2008). This issue is further avoided through obtaining the posterior point estimate of $\mathbf{Z}_{n}$ defined in Equation 8, which minimizes the posterior expectation of a well-chosen loss function. For $\boldsymbol{\sigma}$, or equivalently $\boldsymbol{\phi}$, the likelihood is not invariant to its permutations as it is an ordering. The absence of identifiability issues for $\boldsymbol{\sigma}$ is further evidenced in our applications in Section 3, where the posterior estimates closely approximate the known true topological ordering in the simulated examples (Figure 3), and the posterior distributions of $\boldsymbol{\phi}$ do not look particularly flat for a real dataset (Figure 7).
@@ 3.1 Simulated data
This section presents an illustrative simulation study in which the proposed DAG-SBM (Equation (2)) is compared with the SBM for a directed graph (Equation (1), referred to as the directed SBM). The purpose of this study is to show how fitting the directed SBM without considering the inherent hierarchical ordering among vertices can result in assigning positive posterior probability to many “impossible” edges, and substantial bias in the estimate of the overall number of blocks in the model.
@@
Upon sampling the parameters from their priors specified in Section 2.2, adjacency matrices are sampled from the DAG-SBM for different sample sizes $n=\{250$,500500,1000}1000\} and combinations of the PY parameters, $(\alpha,\theta)=\{(-1,10)$, $(0,1)$, $(0.1,1)$, (0.6,1)}(0.6,1)\}, thus including both finite and infinite regimes. Both the DAG-SBM and directed SBM are fitted to the resulting adjacency matrices, also under both regimes.
@@
Figure 1: Actual adjacency matrix (left) for data simulated from the DAG-SBM (Equation (2)) with $n=250$ and different values of $(\alpha,\theta)$, and estimated probabilities by the directed SBM (Equation (1)) under the finite (middle column) and infinite (right) regime, ordered by true $\boldsymbol{\sigma}$. The middle and right plots correspond to the results in finite and infinite regimes in the corresponding bottom left panels of Figure 2. The dashed line is the major diagonal.
@@
Figure 1 plots the actual adjacency matrix $\mathbf{Y}^{\boldsymbol{\sigma}}$ with $n=250$ (not shown for $n=500$ and $n=1000$), ordered according to the true topological ordering, alongside with the posterior point estimates of $\mathbf{Y}^{\boldsymbol{\sigma}}$, for each simulated dataset fit by the directed SBM. The true adjacency matrices in the first column are necessarily upper triangular, as they are reordered according to the true $\boldsymbol{\sigma}$. However, from the second and third columns, it is evident that many edges in the lower triangular sections have positive posterior probability, while it should be equal to zero. This is especially true for vertices having either high in-degree or out-degree counts. This is due to the fact that the directed SBM is row and column exchangeable, and vertices with high in-degrees tend to also have high out-degree counts. However, when a hierarchical ordering is present, the model is not exchangeable anymore and the out- and in-degrees of each vertex are negatively correlated, with vertices at the beginning of $\boldsymbol{\sigma}$ having higher out-degree than in-degree counts, and the opposite for vertices at the end of $\boldsymbol{\sigma}$.
@@
Figure 2: Posterior histogram of $K_{n}$ for data simulated from the DAG-SBM (Equation (2)), with different values of $(\alpha,\theta,n)$, fitted by both the DAG-SBM (top row) and directed SBM (Equation (1), bottom row), under the finite (solid) and infinite (stripe) regimes. The vertical dashed line is the true value of $K_{n}$.
Figure 2 displays the posterior histograms of the number of blocks $K_{n}$ estimated using the DAG-SBM and the directed SBM, for the same sets of simulated data. On one hand, the directed SBM, which does not include the topological ordering in the likelihood, has the tendency of heavily underestimating or overestimating $K_{n}$. The increase in the number of blocks also negatively affects the computational costs of the algorithm, compared to the DAG-SBM. A comparison of the computational times between the two models is provided in the online Appendix D.1. On the other hand, the DAG-SBM provides reasonable estimates of $K_{n}$ across different sample sizes and for both regimes of the PY parameters. Furthermore, for the DAG-SBM, the model selection step in the MCMC sampler successfully recovers the true regime of the PY parameters. Indeed, Table 1 displays the estimated Bayes factor for the model selection step between the finite and infinite regime, for a range of different values of the hyperparameters of the PY process, including both logarithm (corresponding to the Dirichlet Process, $\alpha=0$) and polynomial growth ($0<\alpha<1$) for $K_{n}$. The estimated Bayes factor suggests that the posterior sampler is capable of recovering the correct regime from the data, particularly when $\alpha$ is far from 0, and being generally indifferent between the regimes when $\alpha$ is close to 0, and $K_{n}$ grows logarithmically, hence very slowly.
$\alpha$ $\theta$ Regime $n=250$ $n=500$ $n=1000$
1-1 $10$ Finite $8.30$ $612.58$ $2.72$
0 11 Infinite: Logarithm growth $2.56$ $0.85$ $19.12$
$0.1$ 11 Infinite: Polynomial growth $1.11$ $0.40$ $0.89$
$0.6$ 11 Infinite: Polynomial growth $7.93\times 10^{-4}$ $7.62\times 10^{-7}$ $1.92\times 10^{-11}$
@@
Table 1: Estimated Bayes factor for the finite regime to the infinite regime when applying the model selection in the MCMC sampler for the DAG-SBM to the simulated data, for different values of $(\alpha,\theta,n)$. A value above (below) 11 favours the finite (infinite) regime.
@@
Finally, we look at the posterior of $\boldsymbol{\sigma}$, or equivalently the positions of the vertices in the topological ordering, $\boldsymbol{\phi}$. The mixing of the MCMC is good, and the trace plots are available for some components of $\boldsymbol{\phi}$ in the online Appendix D. The posterior density of each component of $\boldsymbol{\phi}$ is plotted as a row in Figure 3, with the rows in the true topological ordering. The concentration of posterior mass on the main diagonal suggests that this latent ordering can be well recovered without identifiability issues.
@@
Figure 3: The posterior density of the positions of the vertices in $\boldsymbol{\sigma}$ for $n=250$ and different values of $(\alpha,\theta)$ for data simulated from the DAG-SBM (Equation (2)), fitted by the DAG-SBM under the true regime according to $\alpha$. Each row represents the posterior density of a component of $\boldsymbol{\phi}$.